The 2.5-kg collar shown has a coefficient of kinetic friction ?k=0.225 with the shaft. The spring is unstretched when s=0 and the collar is given an initial velocity of v0=11.4 m/s. The unstretched length of the spring is d=1.5 m and the spring constant is k=98.5 N/m.


The 2.5-kg collar shown has a coefficien
Answer #1


KEi = KEf + The PEf.

Operators to
MU^2 /2 = MV^2 /2 + KX^2 /2
m = 2.5 kg
U = 11.4 m/s

2*11.4^2 = 2*v^2 + 98.5*(1.5)^2

So, v = 4.37 m/s